OpenStudy (anonymous):

NEED IMMEDIATE HELP!! (show work) 1. x/3 + 10 = 15 2. y/2 - 7 =9 3. 3b-4 = 11 4. 12h + 5 = 41 5. -5g + 7 = -38 6. -13r - 14 = -131 7. m/7 - 4 = 2 8. 11k/4- 4 = 18 9. 3f/4 - 9 = 27

OpenStudy (anonymous):

do you not know how to solve any of these problems?

OpenStudy (anonymous):

nah these are bonus questions

OpenStudy (anonymous):

that i dont get

OpenStudy (anonymous):

let me get you started... for the first one we have x/3 + 10 = 15 so we want to subtract 10 from the 15 to leave us with x/3 = 5 and multiply by the reciprocal to get the answer for x

OpenStudy (anonymous):

i forgot what a reciprocal is... (my problem because I came a bit late to math class)

OpenStudy (anonymous):

multiply both sides right?

OpenStudy (anonymous):

the reciprocal is the inverse of a number so x/3 = 1/3 so you would multiply by 3/1

OpenStudy (anonymous):

so its 3/1 * x/3 = 5 * 3/1?